# Superconductivity Proximate to Non-Abelian Fractional Spin Hall Insulator in Twisted Bilayer MoTe2

### v1

@@ Abstract
Twisted bilayer MoTe2 near two-degree twists has emerged as a platform for exotic correlated topological phases, including ferromagnetism and a non-Abelian fractional spin Hall insulator. Here we reveal the unexpected emergence of an intervalley superconducting phase that intervenes between these two states in the half-filled second moiré bands. Using a continuum model and exact diagonalization, we identify superconductivity through multiple signatures: negative binding energy, a dominant pair-density eigenvalue, finite superfluid stiffness, and pairing symmetry consistent with a time-reversal-symmetric nodal extended $s$-wave state. Remarkably, our numerical calculation suggests a continuous transition between superconductivity and the non-Abelian fractional spin Hall insulator, in which topology and symmetry evolve simultaneously, supported by an effective field-theory description. Our results establish higher moiré bands as fertile ground for intertwined superconductivity and topological order, and point to experimentally accessible routes for realizing superconductivity in twisted bilayer MoTe2.
@@ Introduction
Flat topological bands with strong electron correlations have recently attracted significant attention for realizing long-sought topological phases [1, 2, 3, 4, 5], such as fractional Chern insulators [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], which are promising candidates for topological quantum computation [19]. A particularly intriguing platform is small-angle twisted bilayer MoTe2, where the second moiré bands are predicted to host non-Abelian fractional Chern insulators [20, 21, 22, 23, 24, 25, 26], while experiments have uncovered an even more exotic phase—a non-Abelian fractional spin Hall insulator [13, 14]. Notably, in these flat Chern band systems, chiral superconductivity has also been experimentally observed [27], generating considerable interest in uncovering the relationship between superconductivity and topological phases [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39].
@@
In this manuscript, we identify another intriguing quantum state, namely an intervalley superconducting phase, that can arise in this strongly correlated topological flat band system. Specifically, we report evidence for such a phase proximate in the phase diagram to the non-Abelian fractional spin Hall insulator in the half-filled second moiré bands, i.e., hole filling $\nu_{h}=3$, of approximately 2-twisted bilayer MoTe2 [13, 14, 40]. We perform exact diagonalization on a continuum model of the half-filled second moiré bands, incorporating both valleys and screened Coulomb interactions [40]. Superconductivity emerges when the short-range component of the intervalley interactions is screened, yielding time-reversal symmetry and extended $s$-wave pairing. Strikingly, we also find that the superconducting ground state is continuously connected to one of the fractional spin Hall insulator across the phase boundary, closely resembling the continuous transition from bilayer fractional quantum Hall states of two Pfaffian copies to an exciton condensate [41, 42]. Indeed, we can formulate an effective field theory for the unconventional continuous transition between the fractional spin Hall insulator and a superconductor, where topology and symmetry change simultaneously.
@@
Interestingly, when the second moiré bands in our model are replaced by first Landau levels [40], the superconductivity becomes much less apparent. This highlights the importance of realistic band geometry for stabilizing superconductivity, in line with recent studies [28, 29, 30, 31, 32, 33, 34, 35]. Consistent with this picture, we find that a generalized first Landau level and its time-reversal conjugate endowed with non-uniform quantum geometry [43, 44, 45] can reproduce key features of the superconductivity observed in the continuum model of twisted bilayer MoTe2. In summary, our results in this manuscript suggest that higher moiré bands may provide a promising platform to explore the interplay between superconductivity, strong correlations, geometry, and topology, thereby broadening the landscape of superconductivity and fractionalization beyond previous studies [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 46, 47].
@@
Figure 1: Schematic phase diagram and evolution of many-body spectrum. (a) Schematic phase diagram as a function of $d$ in Eq. (2), with $a_{M}=1$. As $d$ increases from 0 to $\infty$, the system evolves through a ferromagnetic Chern insulator, a superconductor, and finally a non-Abelian fractional spin Hall insulator. (b) The lowest energy of each magnetization subspace for $N=14$. (c,d) Evolution of the many-body spectrum within the paramagnetic sector $M_{z}=0$ for $N=14$ and 12, respectively. The evolution of the many-body spectrum suggests there are two transition points, a first-order transition $d_{c,1}$ which changes the magnetization, and another continuous transition at $d_{c,2}$ which changes the nature of the ground state manifold and excitations.
@@ Results
$V_{s_{0}s_{1}}\left(\mathbf{p}\right)=\frac{e^{2}}{4\pi\epsilon}\begin{dcases}\frac{1}{p}&s_{0}=s_{1}\\ \frac{e^{-pd}}{p}&s_{0}\neq s_{1}\end{dcases}.$ (2)
@@
A few remarks are in order. First, at a twist angle near 2, the $K$-valley of twisted bilayer MoTe2 realizes a sequence of Chern bands, all with $C=+1$, closely resembling the sequential Landau levels of a conventional two-dimensional electron gas [51]. The resemblance is not limited to the band topology alone: the effective interactions within each moiré bands strikingly mirror the Haldane pseudopotentials of their corresponding Landau levels [21]. Secondly, the moiré bands possess almost ideal dispersion and geometry [20, 21, 22, 23, 24, 25], such that a half-filled single-valley second moiré band—remarkably akin to the first Landau level—can stabilize a non-Abelian fractional Chern insulator [20, 21, 22, 23, 24, 25, 26], specifically the Pfaffian state [20]. This may be relevant to the resistivity plateau near the filling $\nu_{h}=3/2$ observed in experiment [13, 14]. The band dispersion and quantum geometry of our model can be found in [50].
Figure 2: Many-body spectrum and spectral flow for $N=14$. (a) Many-body spectrum at $d/a_{M}=0.80$ (blue circle and gray x markers). (b) Spectral flow under time-reversal symmetric flux insertion at $d/a_{M}=0.80$. We find a $4\pi$-periodic spectral flow which mixes the four lowest states (blue circles), consistent with the fractional spin Hall insulator. (c) Many-body spectrum at $d/a_{M}=0.28$ (red square, blue x, and gray x markers). Blue x markers are the three other wavefunctions with greatest similarity to the ground state wavefunctions of fractional spin Hall insulator. (d) Under time-reversal symmetric flux insertion of (c), the unique ground state at $d/a_{M}=0.28$ remains isolated from the excited states. The spectral flow is $2\pi$-periodic.
@@
We map out the phase diagram [Fig. 1(a)] of the half-filled second moiré bands, i.e., hole filling $\nu_{h}=3$, as $d$ in Eq. (2) is tuned, which hosts both ferromagnetic Chern insulators and fractional quantum spin Hall insulators, as well as a possible intermediate phase. Notably, both ferromagnetism and fractional quantum spin Hall insulators have already been observed experimentally in different devices at the corresponding filling near 2 twist [13, 14, 52].
@@
To build intuition, we first examine the two limiting cases of Eq. (2), $d\to 0$ and $d\to+\infty$. In the $d\to 0$ limit, intervalley and intravalley interactions become indistinguishable, and Stoner’s mechanism drives ferromagnetism. The holes then spontaneously choose one valley to occupy, giving rise to magnetization—or equivalently, valley polarization. Because the occupied band carries a Chern number, this realizes a ferromagnetic Chern insulator. At the opposite limit, $d\to+\infty$, the valleys decouple entirely. Each half-filled moiré band, forming a time-reversal pair, stabilizes the Pfaffian state and its time-reversal conjugate, together yielding the non-Abelian fractional quantum spin Hall insulator [40]. As both are gapped phases, their stability is expected to extend to finite $d$ [Fig. 1(a)].
@@
We employ exact diagonalization (ED) to systematically examine the stability of both phases with varying $d$. ED is performed with periodic boundary conditions on system sizes $N=10,12,14,$ and $16$ [50]. Our ED results largely confirm the intuition developed above, but also reveal an unexpected intermediate phase [Fig. 1(a)]. For example, the ground-state energies in each magnetization sector [Fig. 1(b)] show that the model stabilizes a ferromagnetic Chern insulator for $d<d_{c,1}\approx 0.15~a_{M}$ as expected, while it becomes paramagnetic for $d>d_{c,1}$. Similarly, we observe clear stabilization of the fractional quantum spin Hall insulator for $d>d_{c,2}\approx 0.33~a_{M}$ [Fig. 1(c,d)], where the system shows a ground-state degeneracy of 36 for $N=12,16$ [50] and 4 for $N=14$ [Fig. 2(a)], consistent with the even–odd effect expected for this phase [53]. The $4\pi$-periodic spectral flow of the ground states in this regime under time-reversal symmetric flux insertion further supports this [Fig. 2(b)].
@@
Beyond the two anticipated phases, we uncover an intermediate phase, characterized by a singly-degenerate paramagnetic ground state appearing for $d\in(d_{c,1},d_{c,2})$ [Fig. 2(c)]. The change in ground-state degeneracy compared to the fractional quantum spin Hall insulator ($d>d_{c,2}$) is not merely a finite-size effect, as evidenced by the distinct evolution of the many-body spectrum under time-reversal symmetric flux insertion. Unlike the $4\pi$-periodic spectral flow of the fractional quantum spin Hall insulator, the unique ground state of the intermediate phase remains isolated from excited states and exhibits a $2\pi$-periodic spectral flow [Fig. 2(d)]. Finally, we also find that the ground-state energy derivatives and ground-state fidelity indicate a continuous phase transition at $d\approx d_{c,2}$ between the intermediate phase and the fractional quantum spin Hall insulator [50]. Both the features and spectral flow of the ground states closely resemble those of the continuous transition from bilayer fractional quantum Hall states of two Pfaffian copies to an exciton condensate [41, 42, 50]. As we will show later, this continuous transition is consistent with our identification of the intermediate state as a superconducting phase.
@@
In [50], we also present the many-body spectra and their evolution in the generalized first Landau level and its time-reversal conjugate with non-uniform quantum geometry. In the idealized first Landau level at $N=14$ with uniform quantum geometry, the system appears to undergo a direct transition from the ferromagnetic Chern insulator to the fractional spin Hall insulator, with no intermediate phase showing a 1-fold degenerate ground state [40]. However, upon introducing non-uniform quantum geometry via a spatially modulated periodic magnetic field [43, 44, 45], the intermediate state emerges and becomes visible. This demonstrates that non-uniform quantum geometry plays a crucial role in stabilizing the intermediate phase, even without explicit band dispersion.
@@
Figure 3: Signatures of superconductivity. (a) Intervalley entanglement entropy, (b) Binding energy $E_{b}^{(\pm)}$, (c) Spectrum of the normalized pair density matrix $\rho_{\mathbf{k}^{\prime}\mathbf{k}}/N$, and (d) Superfluid stiffness $D_{S}$ of the ground state over $d$ for $N=14$. (e.f) Cooper-pair order parameter for $N=12$ and $16$, respectively. The size of the circle is proportional to the absolute value of the order parameter at the corresponding momenta, and red and cyan colors indicate opposite signs.
@@
We now present several pieces of evidence indicating that the intermediate state is a superconductor. Our first clue from the many-body spectrum is that the ground state carries zero total magnetization and momentum across all numerically-accessible system sizes and neighboring fillings in the intermediate $d\in(d_{c,1},d_{c,2})$ [50]. This is consistent with Cooper pairing between holes of opposite spins and valleys at opposite momenta, i.e., between $\left(\uparrow,\mathbf{k}\right)$ and $\left(\downarrow,-\mathbf{k}\right)$. An additional signature is seen in the intervalley entanglement entropy, which changes abruptly from large values in the intermediate regime $d\in(d_{c,1},d_{c,2})$ to small values in the fractional spin Hall insulators ($d>d_{c,2}$) [Fig. 3(a)]. This behavior further supports the picture of intervalley pairing, which inherently links and entangles the two valleys [Fig. 1(a)].
We further observe that the intermediate state exhibits a quantitatively enhanced tendency to form Cooper pairs, as indicated by the binding energy $E_{b}^{(\pm)}=E_{N\pm 2,0}+E_{N,0}-2E_{N\pm 1,0}$, where $E_{N_{h},0}$ is the ground-state energy for $N_{h}$ holes. We find that the binding energy is negative for all $d$, with its magnitude $|E_{b}|$ reaching its maximum in the intermediate region $d\in(d_{c,1},d_{c,2})$ [Fig. 3(b)]. The identification of the intermediate state as a superconductor is further corroborated by the spectrum of the pair density matrix [54],
@@
where $\hat{\psi}_{s,\mathbf{k}}^{\dagger}$ is the operator that creates a hole in the second moiré band with spin $s$ and crystal momentum $\mathbf{k}$. We observe that the pair density matrix shows a single dominant eigenvalue, well separated from the rest of the spectrum, and also this eigenvalue reaches its maximum within the intermediate regime $d\in(d_{c,1},d_{c,2})$ [Fig. 3(c)]. Both results show a tendency for binding, specifically between $\left(\uparrow,\mathbf{k}\right)$ and $\left(\downarrow,-\mathbf{k}\right)$, which is maximized in the intermediate region.
@@
In [50], we report several additional observations. First, as expected, we observe that the many-body spectrum in our model is adiabatically connected to that of an ideal extended $s$-wave superconductor, realized when a strong attractive interaction in this pairing channel is introduced. In addition, we have also computed the charge gap, whose behavior is again consistent with the nodal superconductivity. Finally, analysis of the excitonic pair density matrix rules out intervalley exciton condensation as a candidate for the intermediate phase.
@@
Our ED calculations reveal a direct continuous transition from the non-Abelian fractional quantum spin Hall insulator to an $s$-wave superconductor, where topology and symmetry change simultaneously. Such a transition, lying beyond the traditional Landau–Ginzburg framework of symmetry breaking, warrants further theoretical explanation. We provide a field theoretic description of this continuous transition from the fractional quantum spin Hall insulator to the superconductor, building on the $\mathrm{U}(2)_{2,-4}\times\mathrm{U}(1)_{8}$ Chern-Simons theory of the Pfaffian state [57, 58, 59, 60].
@@
where $a,b$ are the dynamical $\mathrm{U}(2)$ gauge fields and $A$ is the background electromagnetic gauge field. The superconductivity emerges from the condensation of a matter field $\Phi$ in the fundamental representation of both $\mathrm{U}(2)$s, coupled to the gauge fields as $\mathop{}\!\mathrm{D}\Phi=\partial\Phi-ia\Phi-ib\Phi$, which effectively enforces a=ba=-b. This reduces the theory to
@@
which is precisely the effective field theory of the superconductor [28]. In passing, we note that the above construction generalizes straightforwardly to the case of $\mathrm{aPf}\times\overline{\mathrm{aPf}}$, with $\mathrm{aPf}$ denoting the anti-Pfaffian order [50].

### v2

@@ Abstract
Twisted bilayer MoTe2 near two-degree twists has emerged as a platform for exotic correlated topological phases, including ferromagnetism and a non-Abelian fractional spin Hall insulator. Here we reveal the unexpected emergence of an intervalley superconducting phase that intervenes between these two states in the half-filled second moiré bands. Using a continuum model and exact diagonalization, we identify superconductivity through multiple signatures: negative binding energy, a dominant pair-density eigenvalue, finite superfluid stiffness, and pairing symmetry consistent with a time-reversal-symmetric nodal extended $s$-wave state. Remarkably, our numerical calculation suggests a continuous transition between superconductivity and the non-Abelian fractional spin Hall insulator, in which topological order and symmetry evolve simultaneously, supported by an effective field-theory description. Notably, our field-theoretic analysis indicates that superconductivity is driven by the condensation of charge-$e/2$ self-bosonic non-Abelian anyons, thereby providing a concrete realization of anyon superconductivity. Complementarily, when approached from the normal metallic side, superconductivity instead emerges from a Kohn–Luttinger instability enabled by the non-uniform quantum geometry of the flat moiré bands. Our results establish higher moiré bands as fertile ground for intertwined superconductivity and topological order, and point to experimentally accessible routes for realizing superconductivity in twisted bilayer MoTe2.
@@ Introduction
Flat topological bands with strong electron correlations have recently attracted significant attention for realizing long-sought topological phases [12, 14, 11, 68, 10], such as fractional Chern insulators [96, 9, 100, 62, 36, 67, 63, 39, 38, 55, 54, 93, 4], which are promising candidates for topological quantum computation [58]. A particularly intriguing platform is small-angle twisted bilayer MoTe2, where the second moiré bands are predicted to host non-Abelian fractional Chern insulators [3, 66, 13, 94, 83, 52], while experiments have uncovered an even more exotic phase—a non-Abelian fractional spin Hall insulator [39, 38]. Notably, in these flat Chern band systems, chiral superconductivity has also been experimentally observed [97], generating considerable interest in uncovering the relationship between superconductivity and topological phases [24, 95, 79, 80, 65, 86, 103, 59, 78, 30].
@@
In this manuscript, we identify another intriguing quantum state, namely an intervalley superconducting phase, that can arise in this strongly correlated topological flat band system. Specifically, we report evidence for such a phase proximate in the phase diagram to the non-Abelian fractional spin Hall insulator in the half-filled second moiré bands, i.e., hole filling $\nu_{h}=3$, of approximately 2-twisted bilayer MoTe2 [39, 38, 2]. We perform exact diagonalization on a continuum model of the half-filled second moiré bands, incorporating both valleys and screened Coulomb interactions [2]. Superconductivity emerges when the short-range component of the intervalley interactions is screened, yielding time-reversal symmetry and extended $s$-wave pairing.
We further propose two complementary pairing mechanisms underlying the observed superconductivity: a Kohn–Luttinger pairing instability [42] on the metallic side and anyon superconductivity [48, 49] emerging from the non-Abelian spin Hall insulator. The former originates from the screening of repulsive Coulomb interactions, which generates an effective attractive pairing channel, similar to recent studies of superconductivity in the first moiré band of twisted bilayer MoTe2 [24, 95]. Crucially, we find that the non-uniform quantum geometry of the moiré bands plays an essential role in enabling intervalley pairing. For example, when the second moiré bands in our model are replaced by first Landau levels with uniform quantum geometry, the superconducting tendency is strongly suppressed [2]. This comparison underscores the importance of realistic band geometry in stabilizing superconductivity, in agreement with recent theoretical developments [24, 95, 79, 80, 65, 86, 103, 59, 78, 30, 74, 34].
Remarkably, our numerics uncovers a continuous single transition from the fractional spin Hall insulator to the superconducting state, with spectral flows closely resembling those of the continuous transition from a bilayer fractional quantum Hall state composed of two Pfaffian copies to an exciton condensate [75, 105]. This transition is further supported by our effective field-theory description, where superconductivity emerges from the fractional spin Hall insulator via the condensation of non-Abelian charge-$e/2$ anyons. In this description, the transition is unconventional and and lies beyond the traditional Landau–Ginzburg paradigm: it simultaneously alters the global charge-conservation symmetry and the underlying topological order. At the same time, it provides a pairing mechanism in which superconductivity arises from the condensation of charged anyons rather than a Cooper pair, a mechanism that has recently attracted significant attention [79, 80, 65, 86, 103, 59, 78, 30, 101, 44, 77, 73].
@@
In summary, our results in this manuscript suggest that higher moiré bands may provide a promising platform to explore the interplay between superconductivity, strong correlations, quantum geometry, and topological orders, thereby broadening the landscape of superconductivity and fractionalization beyond previous studies [79, 80, 65, 86, 103, 59, 78, 30, 101, 44, 77, 73, 48, 49].
@@
Figure 1: Schematic phase diagram and evolution of many-body spectrum. (a) Schematic phase diagram as a function of $d$ in Eq. (2), with $a_{M}=1$. As $d$ increases from 0 to $\infty$, the system evolves through a ferromagnetic Chern insulator, a superconductor, and finally a fractional spin Hall insulator. (b) The lowest energy of each magnetization subspace for $N=14$. (c,d) Evolution of the many-body spectrum within the paramagnetic sector $M_{z}=0$ for $N=14$ and 12, respectively. The evolution of the many-body spectrum suggests there are two transition points, a first-order transition $d_{c,1}$ which changes the magnetization, and another continuous transition at $d_{c,2}$ which changes the nature of the ground-state manifold and excitations.
@@ Results
$V_{s_{0}s_{1}}\left(\mathbf{p}\right)=\frac{e^{2}}{2\epsilon}\begin{dcases}\frac{1}{p}&s_{0}=s_{1}\\ \frac{e^{-pd}}{p}&s_{0}\neq s_{1}\end{dcases}.$ (2)
@@
A few remarks are in order. First, at a twist angle near 2, the $K$-valley of twisted bilayer MoTe2 realizes a sequence of Chern bands, all with $C=+1$, closely resembling the sequential Landau levels of a conventional two-dimensional electron gas [76]. The resemblance is not limited to the band topology alone: the effective interactions within each moiré bands strikingly mirror the Haldane pseudopotentials of their corresponding Landau levels [3]. Secondly, the moiré bands possess almost ideal dispersion and geometry [3, 66, 13, 94, 83, 52], such that a half-filled single-valley second moiré band—remarkably akin to the first Landau level—can stabilize a non-Abelian fractional Chern insulator [3, 66, 13, 94, 83, 52], specifically the Pfaffian state [66]. This may be relevant to the resistivity plateau near the filling $\nu_{h}=3/2$ observed in experiment [39, 38]. The band dispersion and quantum geometry of our model can be found in [1].
Figure 2: Many-body spectrum and spectral flow for $N=14$. (a) Many-body spectrum at $d/a_{M}=0.80$ (blue circle and gray x markers). (b) Spectral flow under time-reversal symmetric flux insertion at $d/a_{M}=0.80$. We find a $4\pi$-periodic spectral flow which mixes the four lowest states (blue circles), consistent with the fractional spin Hall insulator. (c) Many-body spectrum at $d/a_{M}=0.28$ (red square, magenta, blue, cyan, and gray x markers). The red square (ground-state) and two blue x markers are the wavefunctions with greatest similarity to three of the four ground-state wavefunctions of the fractional spin Hall insulator. The two cyan x markers both have large contributions from the last fractional spin Hall insulator ground-state wavefunction. The first excited state is the magenta x marked state, which is not adiabatically connected to the fractional spin Hall insulator ground-state wavefunctions. (d) Under time-reversal symmetric flux insertion of (c), the unique ground-state at $d/a_{M}=0.28$ remains isolated from the excited states. The spectral flow is $2\pi$-periodic.
@@
We map out the phase diagram [Fig. 1(a)] of the half-filled second moiré bands, i.e., hole filling $\nu_{h}=3$, as $d$ in Eq. (2) is tuned, which hosts both ferromagnetic Chern insulators and fractional spin Hall insulators, as well as a possible intermediate phase. Notably, both ferromagnetism and fractional spin Hall insulators have already been observed experimentally in different devices at the corresponding filling near 2 twist [39, 38, 50].
@@
To build intuition, we first examine the two limiting cases of Eq. (2), $d\to 0$ and $d\to+\infty$. In the $d\to 0$ limit, intervalley and intravalley interactions become indistinguishable, and Stoner’s mechanism drives ferromagnetism. The holes then spontaneously choose one valley to occupy, giving rise to magnetization—or equivalently, valley polarization. Because the occupied band carries a Chern number, this realizes a ferromagnetic Chern insulator. At the opposite limit, $d\to+\infty$, the valleys decouple entirely. Each half-filled moiré band, forming a time-reversal pair, stabilizes the Pfaffian state and its time-reversal conjugate, together yielding the fractional spin Hall insulator [2]. As both are gapped phases, their stability is expected to extend to finite $d$ [Fig. 1(a)].
@@
We employ exact diagonalization (ED) to systematically examine the stability of both phases with varying $d$. ED is performed with periodic boundary conditions on system sizes $N=12,14,$ and $16$ [1]. Our ED results largely confirm the intuition developed above, but also reveal an unexpected intermediate phase [Fig. 1(a)]. For example, the ground-state energies in each magnetization sector [Fig. 1(b)] show that the model stabilizes a ferromagnetic Chern insulator for $d<d_{c,1}\approx 0.15\,a_{M}$ as expected, while it becomes paramagnetic for $d>d_{c,1}$. Similarly, we observe clear stabilization of the fractional spin Hall insulator for $d>d_{c,2}\approx 0.34\,a_{M}$ [Fig. 1(c,d)], where the system shows a ground-state degeneracy of 36 for $N=12,16$ [1] and 4 for $N=14$ [Fig. 2(a)], consistent with the even–odd effect expected for this phase [61]. The $4\pi$-periodic spectral flow of the ground-states in this regime under time-reversal symmetric flux insertion, confirmed by tracking the wavefunction fidelities, further supports this [Fig. 2(b)].
@@
Beyond the two anticipated phases, we uncover an intermediate phase, characterized by a singly-degenerate paramagnetic ground-state appearing for $d\in(d_{c,1},d_{c,2})$ [Fig. 2(c)]. Tracking the wavefunction fidelities, we find that the other low-lying states that previously belonged to the ground-state manifold of the fractional spin Hall insulator now hybridize with excited states. Furthermore, the change in ground-state degeneracy compared to the fractional spin Hall insulator ($d>d_{c,2}$) is not merely a finite-size effect, as evidenced by the distinct evolution of the many-body spectrum under time-reversal symmetric flux insertion. Unlike the $4\pi$-periodic spectral flow of the fractional spin Hall insulator, the unique ground-state of the intermediate phase remains isolated from excited states and exhibits a $2\pi$-periodic spectral flow [Fig. 2(d)]. The change in many-body topology is further evidenced by the spin Hall conductivity which changes from $\sigma_{s}=e^{2}/h$ for $d>d_{c,2}$ to $\sigma_{s}=0$ for $d_{c,1}<d<d_{c,2}$ [1]. Finally, we also find that the ground-state energy derivatives and ground-state fidelity indicate a continuous phase transition at $d=d_{c,2}$ between the intermediate phase and the fractional spin Hall insulator [1].
Both the features and spectral flow of the ground-states closely resemble those of the continuous transition from bilayer fractional quantum Hall states of two Pfaffian copies to an exciton condensate [75, 105]. As we will show later, this continuous transition is consistent with our identification of the intermediate state as a superconducting phase. In [1], we further present data demonstrating the enhanced transition behavior when the short-range repulsion—specifically, the zeroth Haldane pseudopotential [45]—is explicitly suppressed.
@@
In [1], we also present the many-body spectra and their evolution in the generalized first Landau level and its time-reversal conjugate with non-uniform quantum geometry. In the idealized first Landau level at $N=14$ with uniform quantum geometry, the system appears to undergo a direct transition from the ferromagnetic Chern insulator to the fractional spin Hall insulator, with no intermediate phase showing a 1-fold degenerate ground-state [2]. However, upon introducing non-uniform quantum geometry [84, 21, 53], the intermediate state emerges and becomes visible. This demonstrates that non-uniform quantum geometry plays a crucial role in stabilizing the intermediate phase, even without explicit band dispersion.
@@
Figure 3: Signatures of superconductivity. (a) Intervalley entanglement entropy $S_{0}^{(v)}$, (b) Binding energy $E_{b}^{(+)}$, (c) Spectrum of the normalized pair density matrix $\rho_{\mathbf{k}^{\prime}\mathbf{k}}/N$, and (d) Superfluid stiffness $D_{S}$ of the ground-state over $d$.
@@
[…]
We further observe that the intermediate state exhibits a quantitatively enhanced tendency to form Cooper pairs, as indicated by the binding energy,
$E_{b}^{(\pm)}=E_{N\pm 2,0}+E_{N,0}-2E_{N\pm 1,0},$
where $E_{N_{h},0}$ is the ground-state energy for $N_{h}$ holes. Across all three system sizes $N=12,14,16$, we find that the binding energy is negative for all $d$, with its magnitude $|E_{b}|$ comparable to the Fermi energy $E_{F}\approx 1.50\>\mathrm{meV}$ for the intermediate regime $d\in\left(d_{c,1},d_{c,2}\right)$, significantly smaller for $d>d_{c,2}$, and attaining a maximum slope at $d\approx d_{c,2}$ [Fig. 3(b)]. The identification of the intermediate state as a superconductor is further corroborated by the spectrum of the pair density matrix [98],
@@
where $\hat{\psi}_{s,\mathbf{k}}^{\dagger}$ is the operator that creates a hole in the second moiré band with spin $s$ and crystal momentum $\mathbf{k}$. We observe that the pair density matrix shows a single dominant eigenvalue, well separated from the rest of the spectrum, and also this eigenvalue reaches its maximum within the intermediate regime $d\in(d_{c,1},d_{c,2})$ [Fig. 3(c)]. Both results show a tendency for binding, specifically between $\left(\uparrow,\mathbf{k}\right)$ and $\left(\downarrow,-\mathbf{k}\right)$, which is maximized in the intermediate region. In [1], we make a similar analysis of the excitonic pair density matrix, and rule out intervalley exciton condensation as a candidate for the intermediate phase.
@@
Figure 4: Pairing symmetry and Kohn-Luttinger instability. (a,b) Cooper-pair order parameter for $N=12$ and $16$, respectively. The size of the circle is proportional to the absolute value of the order parameter at the corresponding momenta, and red and cyan colors indicate opposite signs. (c) Random-phase-approximation-screened intervalley interaction. (d) Corresponding leading pairing instability. All calculations are done for $d/a_{M}=0.28$.
@@
In [1], we report several additional observations. First, as expected, we observe that the many-body spectrum in our model is adiabatically connected to that of an ideal extended $s$-wave superconductor, realized when a strong attractive interaction in this pairing channel is introduced. In addition, we have also computed the charge gap, whose behavior is again consistent with a nodal superconductor. Finally, we find that the binding energy and Fermi energy are of comparable magnitude and estimate a relatively short superconducting coherence length [1], placing the intermediate superconducting state in the BCS–BEC crossover regime [15], consistent with experimental observations in moiré flat band superconductors of magic-angle twisted bilayer graphene [60, 82], magic-angle twisted trilayer graphene [64, 40], and first moiré band twisted bilayer MoTe2 [97]. We note that our estimate for superconducting coherence length incorporates quantum geometric contributions [32, 16, 28, 33, 81].
Kohn-Luttinger Instability.
The superconductivity observed in our exact diagonalization calculations can be understood in terms of the Kohn–Luttinger mechanism [42]. In particular, the continuum moiré band model in Eq. (1) exhibits both finite dispersion and non-uniform quantum geometry, the latter effectively contributing to dispersion effects when expressed in the hole representation [24, 1]. As detailed in [1], we find that several factors—the presence of multiple Fermi surface pockets [56], the proximity of van Hove singularities to the Fermi surfaces [23, 57], the large quantum metric associated with the first-Landau-level-like character of the band, non-uniform quantum geometry [74], and spin anisotropy—work together to generate substantial screening of the intervalley interaction [Fig. 4(c)].
We find that, once screening within the random-phase approximation is taken into account, the interaction increases with momentum, naturally favoring a sign-changing superconducting order parameter [74]. Solving the linearized Bardeen–Cooper–Schrieffer gap equation, we identify the leading pairing instability [Fig. 4(d)] as an sign-changing extended $s$-wave state in the $A_{1}$ irreducible representation, consistent with the symmetry and overall structure of the gap function obtained from our exact diagonalization calculations.
@@
Our exact diagonalization calculations reveal a direct continuous transition from the non-Abelian fractional spin Hall insulator to an $s$-wave superconductor, where topological order and symmetry, i.e., charge conservation symmetry, change simultaneously. Here, by topological order we refer to the underlying anyon topological data [88, 41, 89, 46, 47, 5, 17]. In this sense, the superconducting phase corresponds to a trivial fermionic topological order [89, 46, 26], whereas the fractional spin Hall insulator realizes the $\mathrm{Pf}\times\overline{\mathrm{Pf}}$ topological order. Such a transition, lying beyond the traditional Landau–Ginzburg framework of symmetry breaking, warrants further theoretical explanation. We provide a field theoretic description of this continuous transition from the non-Abelian fractional spin Hall insulator to the superconductor, building on the $\mathrm{U}(2)_{2,-4}\times\mathrm{U}(1)_{8}$ Chern-Simons theory of the Pfaffian state [20, 72, 51, 27].
@@
where $a,b$ are the dynamical $\mathrm{U}(2)$ gauge fields and $A$ is the background electromagnetic gauge field.
The superconducting phase emerges from the condensation of the bifundamental field $\Phi$, which is coupled to the gauge fields through $\mathop{}\!\mathrm{D}\Phi=\partial\Phi-ia\Phi-i\Phi b$. Notably, we identify the field $\Phi$ with the anyon $\sigma_{\frac{1}{4}}\otimes\sigma_{\frac{1}{4}}^{\mathcal{T}}$ [1], where $\sigma_{\frac{1}{4}}$ denotes the non-Abelian charge-$e/4$ anyon in the Pfaffian theory [8, 80] and $\sigma_{\frac{1}{4}}^{\mathcal{T}}$ represents its time-reversal partner. This anyon $\Phi$ is self-bosonic, yet exhibits nontrivial braiding statistics with other anyons in the theory, e.g., $\psi_{0}\otimes\sigma_{\frac{1}{4}}^{\mathcal{T}}$, and possesses multiple fusion channels, i.e., it is non-Abelian [1]. To model the condensation transition, we introduce the potential
@@
with the real coefficients $m^{2}$, $\lambda_{1}$, and $\lambda_{2}$. When $m^{2}<0$, the potential $V(\Phi^{\dagger}\Phi)$ is minimized at $\langle\Phi\rangle=sI$ with $s>0$. This condensation enforces the constraint $\alpha\equiv a=-b$ at low energies, leading to the effective theory
[…]
which is precisely the effective field theory of a superconductor [26]. Several technical field-theoretic details of the condensed phase, including the fate of the $\mathrm{U}(2)$ gauge fields, the compactification of $\mathrm{Tr}\,\alpha$, and the absence of gapless Goldstone modes associated with $\Phi$, can be found in [1].
We conclude with a few remarks. First, transitions of this type, known as Chern–Simons–Landau–Ginzburg transitions, have been extensively studied in the literature and are widely believed to be continuous [87, 43, 7, 6, 22, 99, 91, 71, 31, 25, 104, 29, 85, 35]. Our exact diagonalization results are consistent with these observations. Second, the above field-theoretic analysis indicates that the transition is driven by the condensation of the charge-$e/2$ self-bosonic non-Abelian anyon $\Phi$, identified as the composite of $\sigma_{\frac{1}{4}}$ in $\mathrm{Pf}$ and its time-reversal partner $\sigma_{\frac{1}{4}}^{\mathcal{T}}$ in $\overline{\mathrm{Pf}}$. This interpretation is further reinforced by the category-theoretic framework of anyon superconductivity recently developed in [73], whose application to the present system is presented in detail in [1]. Finally, the above construction generalizes straightforwardly to the case of $\mathrm{aPf}\times\overline{\mathrm{aPf}}$, with $\mathrm{aPf}$ denoting the anti-Pfaffian order [1].